# SemanticSplat: Feed-Forward 3D Scene Understanding with Language-Aware Gaussian Fields

### v1

@@ 3.6. Training Loss
(8) $\mathcal{L}_{\text{rgb}}=\sum_{i}\|\mathcal{R_{C}}(\mu,\alpha,\Sigma,c,f)_{i}-I^{\text{gt}}_{i}\|_{1}+\lambda_{1}\cdot\text{LPIPS}(\mathcal{R}_{i},I^{\text{gt}}_{i})$
where $\mathcal{R_{C}}$ is the differentiable renderer of image and $I^{\text{gt}}$ the target image. And the loss weights of LPIPS (Zhang et al., 2018) loss weight $\lambda_{1}$ is set to 0.05.
@@
(9) $\mathcal{L}_{\text{sam}}=L_{dist}^{Seg}+\lambda_{mask}L_{mask}^{Seg}$
@@
(10) $\mathcal{L}_{\text{clip}}=L_{dist}^{Lang}$
4. Experiments
4.1. Settings
@@
We utilize the ScanNet (Dai et al., 2017) dataset for training, which provides high-fidelity 3D geometry and high-resolution RGB images, along with estimated camera intrinsic and extrinsic parameters for each frame. A total of 1,462 scenes are used for training, while 50 unseen scenes are reserved for validation. All frames are cropped and resized to a resolution of 256 × 256.
@@
To evaluate photometric fidelity, we adopt standard image quality metrics: pixel-level PSNR, patch-level SSIM (Wang et al., 2004), and feature-level LPIPS (Zhang et al., 2018). For semantic segmentation, we measure performance using mean Intersection-over-Union (mIoU) and mean pixel accuracy (mAcc).
@@
Our model is trained using Adam (Kingma, 2014) optimizer with an initial learning rate of $41e-$ 4 and cosine decay following. Both semantic fea- ture distillation stages are trained on 4 Nvidia A100 GPU for 5000 iterations.
@@
GT
Ours
LSM
Feature-3DGS-LSeg
Feature-3DGS-SAM
MVSplat
Figure 3. Novel View Synthesis Comparisons. Our method outperforms LSM and Feature-3DGS in challenging regions and is compatible with baseline MVSplat, which shows we reconstruct the appearance successfully
4.2. Holistic Semantic Field Reconstruction
@@
Figure 4. Language-based Segmentation Comparison. We visualize the segmentation from a set of categories for unseen view, our method outperforms with the other 3D method and comparably to the 2D VFMs, which indicates we effictively lift 2d foundation language-image model to 3D.
Table 1. Comparison (Language-Seg). Performance metrics for source and target view segmentation across different methods.
@@
Table 2. Comparison (Promtable-Seg). Performance metrics for source and target view segmentation across different methods.
@@
We compare our approach with two state-of-the-art methods: LSM (Fan et al., 2024) (a generalizable framework) and Feature-3DGS (Zhou et al., 2024b) (a per-scene optimization-based method). Both methods predict RGB values and leverage feature-based 3D Gaussian Splatting (3D-GS) (Kerbl et al., 2023). Unlike our approach, Feature-3DGS supports promptable segmentation and open-vocabulary segmentation by separately optimizing SAM and LSeg features, while LSM is limited to open-vocabulary segmentation.
@@
We further compare our method with the baseline MVSplat (Chen et al., 2024c), a feed-forward Gaussian reconstruction model trained on the RealEstate10K (Zhou et al., 2018) dataset. As shown in Table 1 and Figure 3, Feature-3DGS (Zhou et al., 2024b) struggles to synthesize high-quality images from sparse input views, while LSM (Fan et al., 2024) introduces noisy artifacts, especially near object boundaries. Our method achieves comparable pixel-level quality to MVSplat despite training on lower-quality data and outperforms it at the patch and feature levels, owing to the semantic priors integrated into our pipeline.
@@
Figure 5. Visualization of the Semantic Feature Field. We visualize the language features and segmentation characteristics of the novel views, demonstrating how we elevate the 2D features into 3D while maintaining consistency across views. The visualizations are generated using PCA (Pedregosa et al., 2011).
@@
Following Feature-3DGS (Zhou et al., 2024b), we map thousands of category labels from diverse datasets into a unified set of common categories: {Wall, Floor, Ceiling, Chair, Table, Bed, Sofa, Others}. For comparison, we also include the 2D open-vocabulary segmentation method LSeg (Li et al., 2022). As shown in Table 1 and visualised in Figure 4, our method achieves competitive performance against baseline 3D methods and matches the accuracy of 2D methods when evaluated on the ScanNet dataset with transferred labels. Notably, while LSeg suffers from cross-view inconsistency, our approach maintains high consistency across views. To illustrate this, we visualize the language feature fields of both methods using PCA (projecting high-dimensional features into three channels) (Halko et al., 2011) in Figure 5.
Table 3. Feature‑Condition Ablation (Stages 1 & 2). Stage 1: feature branch under LSeg vs. GT masks; Stage 2: feature branch under SAM vs. GT masks.
@@
Figure 6. Prompt-based Segmentation Comparison. We visualize the segmentation generated from a point prompt for unseen views. Our method outperforms other 3D approaches, indicating that we effectively extend a 2D foundation segmentation model to 3D.
@@
Building on the SAM mask decoder, our method predicts three hierarchical masks from point queries and the predicted segmentation feature map. We uniformly sample a grid of points on the images (32 points along the width and 32 along the height, totaling 1,024 points). For each point, we generate hierarchical masks and evaluate their alignment with ground truth masks from ScanNet labels by reporting the highest Intersection-over-Union (IoU) and accuracy (Acc) scores in Table 2 and visualize the promptable segmentation in Figure 6. In addition to Feature-3DGS, we include the 2D promptable segmenter SAM in our comparisons. To visualize the segmentation feature field and SAM features, we project them into three channels using PCA (Halko et al., 2011) in Figure 5.
4.3. Ablation Studies
@@
Figure 7. Ablation study on different Conditions, HCAM and Mask loss. We visualize the segmentation results under different conditions, illustrating that all these are complementary.

### v2

@@ 3.6 Training Loss
$\mathcal{L}_{\text{rgb}}=\sum_{i}\|\mathcal{R_{C}}(\mu,\alpha,\Sigma,c,f)_{i}-I^{\text{gt}}_{i}\|_{1}+\lambda_{1}\cdot\text{LPIPS}(\mathcal{R}_{i},I^{\text{gt}}_{i})$ (8)
where $\mathcal{R_{C}}$ is the differentiable renderer of image and $I^{\text{gt}}$ the target image. And the loss weights of LPIPS [50] loss weight $\lambda_{1}$ is set to 0.05.
@@
$\mathcal{L}_{\text{sam}}=L_{dist}^{Seg}+\lambda_{mask}L_{mask}^{Seg}$ (9)
@@
$\mathcal{L}_{\text{clip}}=L_{dist}^{Lang}$ (10)
4 Experiments
4.1 Settings
@@
We utilize the ScanNet [9] dataset for training, which provides high-fidelity 3D geometry and high-resolution RGB images, along with estimated camera intrinsic and extrinsic parameters for each frame. A total of 1,462 scenes are used for training, while 50 unseen scenes are reserved for validation. All frames are cropped and resized to a resolution of 256 × 256.
@@
To evaluate photometric fidelity, we adopt standard image quality metrics: pixel-level PSNR, patch-level SSIM [43], and feature-level LPIPS [50]. For semantic segmentation, we measure performance using mean Intersection-over-Union (mIoU) and mean pixel accuracy (mAcc).
@@
Our model is trained using Adam [19] optimizer with an initial learning rate of $41e-$ 4 and cosine decay following. Both semantic fea- ture distillation stages are trained on 4 Nvidia A100 GPU for 5000 iterations.
@@
GTg
Oursg
LSMg
Feat-3DGS-LSeg
Feat-3DGS-SAMg
MVSplatg
Figure 3: Novel View Synthesis Comparisons. Our method outperforms LSM and Feature-3DGS in challenging regions and is compatible with baseline MVSplat, which shows we reconstruct the appearance successfully
4.2 Holistic Semantic Field Reconstruction
@@
Figure 4: Language-based Segmentation Comparison. We visualize the segmentation from a set of categories for unseen view, our method outperforms with the other 3D method and comparably to the 2D VFMs, which indicates we effictively lift 2d foundation language-image model to 3D.
Table 1: Comparison (Language-Seg). Performance metrics for source and target view segmentation across different methods.
@@
Table 2: Comparison (Promtable-Seg). Performance metrics for source and target view segmentation across different methods.
@@
We compare our approach with two state-of-the-art methods: LSM [12] (a generalizable framework) and Feature-3DGS [52] (a per-scene optimization-based method). Both methods predict RGB values and leverage feature-based 3D Gaussian Splatting (3D-GS) [15]. Unlike our approach, Feature-3DGS supports promptable segmentation and open-vocabulary segmentation by separately optimizing SAM and LSeg features, while LSM is limited to open-vocabulary segmentation.
@@
We further compare our method with the baseline MVSplat [6], a feed-forward Gaussian reconstruction model trained on the RealEstate10K [55] dataset. As shown in Table 1 and Figure 3, Feature-3DGS [52] struggles to synthesize high-quality images from sparse input views, while LSM [12] introduces noisy artifacts, especially near object boundaries. Our method achieves comparable pixel-level quality to MVSplat despite training on lower-quality data and outperforms it at the patch and feature levels, owing to the semantic priors integrated into our pipeline.
@@
Figure 5: Visualization of the Semantic Feature Field. We visualize the language features and segmentation characteristics of the novel views, demonstrating how we elevate the 2D features into 3D while maintaining consistency across views. The visualizations are generated using PCA [31].
@@
Following Feature-3DGS [52], we map thousands of category labels from diverse datasets into a unified set of common categories: {Wall, Floor, Ceiling, Chair, Table, Bed, Sofa, Others}. For comparison, we also include the 2D open-vocabulary segmentation method LSeg [22]. As shown in Table 1 and visualised in Figure 4, our method achieves competitive performance against baseline 3D methods and matches the accuracy of 2D methods when evaluated on the ScanNet dataset with transferred labels. Notably, while LSeg suffers from cross-view inconsistency, our approach maintains high consistency across views. To illustrate this, we visualize the language feature fields of both methods using PCA (projecting high-dimensional features into three channels) [14] in Figure 5.
Table 3: Feature‑Condition Ablation (Stages 1 & 2). Stage 1: feature branch under LSeg vs. GT masks; Stage 2: feature branch under SAM vs. GT masks.
@@
Figure 6: Prompt-based Segmentation Comparison. We visualize the segmentation generated from a point prompt for unseen views. Our method outperforms other 3D approaches, indicating that we effectively extend a 2D foundation segmentation model to 3D.
@@
Building on the SAM mask decoder, our method predicts three hierarchical masks from point queries and the predicted segmentation feature map. We uniformly sample a grid of points on the images (32 points along the width and 32 along the height, totaling 1,024 points). For each point, we generate hierarchical masks and evaluate their alignment with ground truth masks from ScanNet labels by reporting the highest Intersection-over-Union (IoU) and accuracy (Acc) scores in Table 2 and visualize the promptable segmentation in Figure 6. In addition to Feature-3DGS, we include the 2D promptable segmenter SAM in our comparisons. To visualize the segmentation feature field and SAM features, we project them into three channels using PCA [14] in Figure 5.
4.3 Ablation Studies
@@
Figure 7: Ablation study on different Conditions, HCAM and Mask loss. We visualize the segmentation results under different conditions, illustrating that all these are complementary.
@@ Conclusion
Appendix A Ablation on Mask Loss
We compare our first-stage feature distillation (full) with a variant that excludes the mask loss (w/o mask loss). We evaluate the segmentation metrics using SAM masks to assess the effectiveness of the distillation process in Table 4.
Table 4: Ablation Study. Impact of mask loss on segmentation.
[…]
w/ mask loss 0.668 0.847
w/o mask loss 0.659 0.842
Appendix B Module Timing
We evaluate the computational cost of each module by running inference on the ScanNet dataset and calculating the runtime for each component of our method, as detailed in Table 5.
Table 5: Inference Time per Module.
Module Time (s)
Depth map predict w/ condition (3.1+3.2) 0.092
Segmentation field distillation (stage 1) 0.032
Language field distillation (stage 2) 0.029
Total 0.153